### v3

# A Class of Generalized Shift-Splitting Preconditioners for Double Saddle Point Problems

Abstract
In this paper, we propose a generalized shift-splitting (GSS) preconditioner, along with its two relaxed variants to solve the double saddle point problem (DSPP). The convergence of the associated GSS iterative method is analyzed, and sufficient conditions for its convergence are established. Spectral analyses are performed to derive sharp bounds for the eigenvalues of the preconditioned matrices. Numerical experiments based on examples arising from the PDE-constrained optimization problems demonstrate the effectiveness and robustness of the proposed preconditioners compared with existing state-of-the-art preconditioners.
Keywords. Double saddle point problem, Preconditioner, GMRES, Shift-splitting, Krylov subspace methods, PDE-constrained optimization.
AMS subject classification. 65F08 $\cdot$ 65F10 $\cdot$ 65F50
1 Introduction
Suppose $n,m$ and $l$ are given positive integers with $n\geq m\geq l.$ Then, we consider the double saddle point problem (DSPP) in the following form [13]:
$\mathcal{B}{\bm{u}}:=\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&D&C\\ -B&-C^{T}&\bm{0}\end{matrix}\right]\left[\begin{matrix}x\\ y\\ z\end{matrix}\right]=\left[\begin{matrix}p\\ q\\ r\end{matrix}\right]=:{\bm{b}},$ (1.1)
where $A\in{\mathbb{R}}^{n\times n}$ is symmetric positive definite (SPD) matrix, $B\in{\mathbb{R}}^{m\times n}$ has full row rank, $C\in{\mathbb{R}}^{l\times m}$ and $D\in{\mathbb{R}}^{l\times l}$ is a symmetric positive semidefinite (SPS) matrix. Further, $p\in{\mathbb{R}}^{n},$ $q\in{\mathbb{R}}^{l}$ and $r\in{\mathbb{R}}^{m}$ are known vectors and $x\in{\mathbb{R}}^{n},$ $y\in{\mathbb{R}}^{l}$ and $z\in{\mathbb{R}}^{m}$ are unknown vectors to be determined. Here, $(\cdot)^{T}$ and $\bm{0}$ represent the transpose of a matrix and zero matrix of appropriate size, respectively.
The DSPP (1.1) is frequently encountered in a wide range of scientific and computational disciplines. Notable areas of application include quadratic programming problems [21], computational fluid dynamics [17], equality constrained indefinite least squares problems[12], PDE-constrained optimization problem [33], and so on.
Owing to the broad applicability of the DSPP (1.1), this paper primarily concentrates on its numerical solution. Nevertheless, the large and sparse nature of the saddle point matrix $\mathcal{B}$ makes the iterative methods more favorable than the direct approaches [35]. In particular, Krylov subspace methods, such as the generalized minimal residual (GMRES) method [36], have emerged prominently as one of the most effective and widely used iterative solvers. However, the ill-conditioned nature and poor spectral properties of the coefficient matrix $\mathcal{B}$ lead to a significant slowdown of the Krylov subspace iterative method. This challenge underscores the need for the development of novel, effective, and robust preconditioners to enhance the performance of Krylov subspace iterative methods.
Assuming $\mathbf{A}=\left[\begin{matrix}A&\bm{0}\\ \bm{0}&D\end{matrix}\right],$ $\mathbf{B}=[B\leavevmode\nobreak\ \leavevmode\nobreak\ C^{T}],$ ${\bf f}=[p^{T},q^{T}]^{T},$ ${\bf g}=r,$ ${\bf x}=[x^{T},y^{T}]^{T}$ and ${\bf y}=z,$ the DSPP can be transformed into the two-by-two block saddle point problems [10]. A surge amount of work has been done in recent years to solve two-by-two block saddle point problems, for instance, Uzawa iterative methods [4, 26, 39], successive overrelaxation methods [6, 19], Hermitian and skew-Hermitian methods [3, 5], and so on. However, these methods do not exploit the three-by-three block structure of $\mathcal{B}.$
To leverage the full block structure of $\mathcal{B},$ various preconditioners have been studied in the literature to solve the DSPP (1.1). When $D=\bm{0}$ in (1.1), block diagonal (BD) and tridiagonal type preconditioners [23, 1], shift-splitting (SS) type preconditioners [14], Uzawa methods [24, 22], etc. have been explored. When $D\neq\bm{0},$ BD preconditioners for the DSPP (1.1) have been investigated in [13]. However, the involvement of the Schur complement matrix makes the computation of the BD preconditioner time-consuming and expensive. By splitting the coefficient matrix $\mathcal{B}$ as $\mathcal{B}=\mathcal{B}_{1}+\mathcal{B}_{2},$ where
$\mathcal{B}_{1}=\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&\bm{0}&\bm{0}\\ -B&\bm{0}&\bm{0}\end{matrix}\right]\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{B}_{2}=\left[\begin{matrix}\bm{0}&\bm{0}&\bm{0}\\ \bm{0}&D&C\\ \bm{0}&-C^{T}&\bm{0}\end{matrix}\right],$ (1.2)
Benzi and Guo [8] proposed the dimensional spitting (DS) preconditioner $\mathcal{P}_{\mathrm{DS}}$ and a relaxed dimensional factorization (RDF) preconditioner $\mathcal{P}_{\mathrm{RDF}}$ [11]. These are given as follows:
$\mathcal{P}_{\mathrm{DS}}=\frac{1}{\alpha}\left[\begin{matrix}\alpha I+A&\bm{0}&B^{T}\\ \bm{0}&\alpha I&\bm{0}\\ -B&\bm{0}&\alpha I\end{matrix}\right]\left[\begin{matrix}\alpha I&\bm{0}&\bm{0}\\ \bm{0}&\alpha I+D&C\\ \bm{0}&-C^{T}&\alpha I\end{matrix}\right],$ (1.3)
$\mathcal{P}_{\mathrm{RDF}}=\frac{1}{\alpha}\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&\alpha I&\bm{0}\\ -B&\bm{0}&\alpha I\end{matrix}\right]\left[\begin{matrix}\alpha I&\bm{0}&\bm{0}\\ \bm{0}&D&C\\ \bm{0}&-C^{T}&\alpha I\end{matrix}\right],$ (1.4)
where $I$ stands for the identity matrix of appropriate size. For more on DS-based preconditioners, refer to [20, 38].
For the DSPP (1.1) arising from PDE-constrained optimization problem, Rees et al. [34] introduced BD preconditioner $\mathcal{P}_{\text{D}}$ and constrained preconditioner $\mathcal{P}_{\text{C}}.$ Later, a block triangular (BT) preconditioner is developed in [32]. In [40], the authors introduced two types of preconditioners for solving the DSPP: a block-counter-diagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCD}}$, and a block-counter-tridiagonal preconditioner, denoted as $\mathcal{P}_{\mathrm{BCT}}$. By using different approximations of Schur complement, preconditioners in BD and BT formats are constructed in [28, 29, 30]. For more research on solving DSPP in the context of PDE-constrained optimization problems, refer to the works [27, 18, 25].
For non-Hermitian linear systems, Bai et al. [7] introduced the SS preconditioner by applying the matrix SS technique to the coefficient matrix. Building on this concept, several SS-type preconditioners have been studied for two-by-two and three-by-three block saddle point problems; see, for example, [14, 15, 16, 37, 2]. However, SS-type preconditioners remain unexplored for DSPP, despite their potential for high efficiency. Motivated by this, in this paper, we introduce a generalized SS (GSS) iterative method and the associated GSS preconditioner for solving the DSPP by implementing the concept of SS to $\mathcal{B}$. Moreover, we derive sufficient criteria for the convergence of the GSS iterative method. The main contributions of the paper are summarized as follows:
  • A novel GSS iterative method and corresponding GSS preconditioner are introduced by implementing the SS approach for the coefficient matrix $\mathcal{B}$ to solve DSPP (1.1).
  • Convergence analysis of the proposed GSS iterative method are carried out, yielding sufficient conditions for its convergence.
  • To further enhance the effectiveness of the GSS preconditioner, two relaxed variants, termed RGSS-I and RGSS-II are proposed, and the spectral bounds of the RGSS-I and RGSS-II preconditioned matrices are thoroughly investigated.
  • Finally, numerical experiments are conducted for the DSPP arising from the PDE-constrained optimization problem to demonstrate the effectiveness of the proposed preconditioners.
The outline of the rest of the paper is as follows. Section 2 investigates the solvability conditions of the DSPP (1.1) and properties of the coefficient matrix $\mathcal{B}.$ The GSS iterative process and associated preconditioner are proposed in Section 3. Section 4 investigates the convergence criteria for the proposed GSS iterative method. Two relaxed variants of the proposed GSS preconditioner are presented in Section 5, and the spectral analyses of the corresponding preconditioned matrices are performed. Section 6 deals with the parameter selection strategies of the proposed preconditioners. Experimental results, analyses and discussions of the proposed and existing preconditioners are discussed in Section 7. At the end, Section 8 includes some concluding statements.
Notation. Throughout the paper, we adopt the following notations. Let $\mathbb{C}^{n}$ and ${\mathbb{R}}^{m\times n}$ denote the collection of all $n$-dimensional complex vectors and $m\times n$ real matrices, respectively. The notations $\|A\|_{F},$ $\sigma(A)$ and $tr(A)$ denote the Frobenious norm, spectrum and trace of the matrix $A\in{\mathbb{R}}^{m\times n},$ respectively. For any vector $x\in\mathbb{C}^{n},$ ${x}^{*}$ represents its conjugate transpose. For any matrix $A\in{\mathbb{R}}^{m\times n}$ having real eigenvalues, $\lambda_{\max}(A)$ and $\lambda_{\min}(A)$ denote the maximum eigenvalue and minimum eigenvalue of $A,$ respectively. For any $z\in\mathbb{C},$ $\mathfrak{R}(z)$ and $\Im(z)$ represent the real and imaginary part of $z,$ respectively.
2 Solvability conditions for the DSPP and properties of the matrix $\mathcal{B}$
In this section, we provide the solvability conditions on the block matrices $A,$ $B,$ $C$ and $D$ for the system (1.1) and a few important properties of the matrix $\mathcal{B}.$
Proposition 2.1.
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be $SPD$ matrices. If $B$ has full row rank, then the saddle point matrix $\mathcal{B}$ is nonsingular.
Proof.
Let $B$ has full row rank, and ${\bm{u}}=[x^{T},y^{T},z^{T}]^{T}\in{\mathbb{R}}^{n+l+m}$ be such that $\mathcal{B}{\bm{u}}=0.$ Then, we have
$\left\{\begin{array}[]{cc}Ax+B^{T}z=&\bm{0},\\ Dy+Cz=&\bm{0},\\ -Bx-C^{T}y=&\bm{0}.\end{array}\right.$ (2.4)
We first assert that $x=\bm{0}.$ Then from the first equation in (2.4), we obtain $B^{T}z=\bm{0}.$ Since, $B$ has full row rank, this implies $z=\bm{0}.$ Thus the second equation of (2.4) gives $y=\bm{0}$ as $D$ is SPD. This implies ${\bm{u}}=\bm{0}.$ Next, we assert $z=\bm{0}.$ Then from the first and second equation of (2.4), we find that $x=\bm{0}$ and $y=\bm{0},$ respectively, as $A$ and $D$ are SPD. Hence, ${\bm{u}}=\bm{0}.$ Now, we assume that $x\neq\bm{0}$ and $z\neq\bm{0}.$ Then multiplying by $x^{T}$ from the left side of the first equation of (2.4), we obtain
$x^{T}Ax+x^{T}B^{T}z=0.$ (2.5)
Again, multiplying third equation of (2.4) by $z^{T}$ from the left, we get
$-z^{T}Bx-z^{T}C^{T}y=0.$ (2.6)
Substituting (2.5) and $Dy=-Cz$ on (2.6), we have $x^{T}Ax+y^{T}Dy=0,$ which is possible only when $x=\bm{0}$ and $y=\bm{0},$ as $A$ and $D$ are SPD matrices. Thus ${\bm{u}}=\bm{0},$ and hence, $\mathcal{B}$ is nonsingular. $\blacksquare$
Proposition 2.2.
Let $A\in{\mathbb{R}}^{n\times n}$ be SPD and $D\in{\mathbb{R}}^{l\times l}$ be SPS. If $B$ and $C$ are of full row rank, then $\mathcal{B}$ is nonsingular.
Proof.
The proof follows similarly to the proof of proposition 2.1.
Proposition 2.3.
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ be SPD matrices. If $B$ and $C$ are of full row rank, then the matrix $\mathcal{B}$ is positive stable, i.e., $\lambda>0$ for all $\lambda\in\sigma(\mathcal{B}).$
Proof.
Suppose $\lambda$ is an eigenvalue of $\mathcal{B}$ and ${\bm{u}}=[u^{T},v^{T},w^{T}]^{T}\in{\mathbb{R}}^{n+l+m}$ is the corresponding eigenvector. Then, we have $\mathcal{B}{\bm{u}}=\lambda{\bm{u}},$ which leads to following three linear system of equations:
$\left\{\begin{array}[]{cc}Au+B^{T}w&=\lambda u,\\ Dv+Cw&=\lambda v,\\ -Bu-C^{T}v&=\lambda w.\end{array}\right.$ (2.10)
Premultiplying $\mathcal{B}{\bm{u}}=\lambda{\bm{u}}$ by ${\bm{u}}^{*},$ we get
$\lambda\|{\bm{u}}\|^{2}=u^{\star}Au+v^{*}Dv+2{i\mkern 1.0mu}\Im(u^{*}B^{T}w+w^{*}Cv).$ (2.11)
Therefore, from (2.11), we obtain
$\mathfrak{R}(\lambda)=\frac{u^{*}Au+v^{*}Dv}{\|{\bm{w}}\|^{2}}.$ (2.12)
From (2.12), we get $\mathfrak{R}(\lambda)\geq 0$ and $\mathfrak{R}(\lambda)=0$ if and only if $u=\bm{0}$ and $v=\bm{0}.$
Next, we show that the vectors $u$ and $v$ can not be zero simultaneously. First, assume that $u=\bm{0}.$ From the first equation in (2.10), it follows that $B^{T}v=\bm{0}$, which leads to $v=\bm{0},$ as $B$ has full row rank. Substituting $u=\bm{0}$ and $v=\bm{0}$ into the third equation of (2.10) yields $u=\bm{0}$. As a result, we obtain $\bm{u}=\mathbf{0}$, which contradicts the assumption that $\bm{u}$ is an eigenvector. Therefore, we conclude that $u\neq\bm{0}$, which in turn implies that $\mathfrak{R}(\lambda)>0$. This completes the proof. $\blacksquare$
3 Proposed generalized shift-splitting (GSS) iterative method and preconditioner
This section proposes a generalized shift-splitting (GSS) iterative method to solve the DSPP. Let $\alpha,\beta,\tau,$ $\bm{\omega}$ be positive real numbers and $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l},$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices, then the saddle point $\mathcal{B}$ admits the following matrix splitting:
$\mathcal{B}=(\Theta+{\bm{\omega}}\mathcal{B})-(\Theta-(1-\bm{\omega})\mathcal{B})=:\mathcal{M}_{\mathrm{GSS}}-\mathcal{N}_{\mathrm{GSS}},$ (3.1)
where
$\mathcal{M}_{\mathrm{GSS}}=\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right],$ (3.2)
$\mathcal{N}_{\mathrm{GSS}}=\left[\begin{matrix}\alpha P-(1-\bm{\omega})A&\bm{0}&(1-\bm{\omega})B^{T}\\ \bm{0}&\beta Q-(1-\bm{\omega})D&(1-\bm{\omega})C\\ -(1-\bm{\omega})B&-(1-\bm{\omega})C^{T}&\tau R\end{matrix}\right],$ (3.3)
and $\Theta=\left[\begin{matrix}\alpha P&\bm{0}&\bm{0}\\ \bm{0}&\beta Q&\bm{0}\\ \bm{0}&\bm{0}&\tau R\end{matrix}\right].$
The special matrix splitting in equation (3.1) introduces a novel iteration scheme, known as the GSS iterative method, for solving the DSPP.
Method 3.1.
(GSS iterative method). Given the initial guess vector ${\bm{u}}^{(0)}=\left[{x^{(0)}}^{T},{y^{(0)}}^{T},{z^{(0)}}^{T}\right]^{T},$ positive real numbers $\alpha,\beta,\tau$ and $\bm{\omega},$ SPD matrices $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m},$ until the stopping criterion is satisfied, compute
${\bm{u}}^{(k+1)}=\mathcal{G}{\bm{u}}^{(k)}+{\bm{d}},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k=0,1,2,\ldots,$ (3.4)
where ${\bm{u}}^{(k)}=\left[{x^{(k)}}^{T},{y^{(k)}}^{T},{z^{(k)}}^{T}\right]^{T}\in{\mathbb{R}}^{n+l+m},$ $\mathcal{G}=\mathcal{M}^{-1}_{\mathrm{GSS}}\mathcal{N}_{\mathrm{GSS}}$ is the iteration matrix and ${\bm{d}}=\mathcal{M}_{\mathrm{GSS}}^{-1}{\bm{b}}\in{\mathbb{R}}^{n+m+l}.$
The matrix splitting in (3.1) induces a preconditioner, denoted as $\mathcal{M}_{\mathrm{GSS}}$, which can be utilized to speed up the convergence rate of the Krylov subspace iterative methods, like GMRES. This preconditioner is referred to as the GSS preconditioner.
At each step of the GSS iterative method or GSS preconditioned GMRES (PGMRES) method, we are required to solve a system of linear equations in the following form:
$\mathcal{M}_{\mathrm{\mathrm{GSS}}}{\bm{z}}=r,$ (3.5)
[…]
$\mathcal{M}_{\mathrm{GSS}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -\bm{\omega}B(\alpha P+\bm{\omega}A)^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\alpha P+\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\widehat{R}\end{matrix}\right]$
.$\hskip 113.81102pt.\left[\begin{matrix}I&\bm{0}&\bm{\omega}(\alpha P+\bm{\omega}A)^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$
where $\widehat{R}=\tau R+\bm{\omega}^{2}B(\alpha P+\bm{\omega}A)^{-1}B^{T}+\bm{\omega}^{2}C^{T}(\beta Q+\bm{\omega}D)^{-1}C.$ In the following, we present the algorithmic implementations of the GSS preconditioner designed to accelerate the GMRES method.
Algorithm 1 Solving $\mathcal{M}_{\mathrm{GSS}}{\bm{z}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\alpha,\beta,\tau,\bm{\omega},$ SPD matrices $P\in{\mathbb{R}}^{m\times m},$ $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
Output: Solution vector ${\bm{z}}=[{\bm{z}}_{1}^{T},{\bm{z}}_{2}^{T},{\bm{z}}_{3}^{T}]^{T}\in{\mathbb{R}}^{n+l+m}.$
Steps:
$1:$ Solve $(\alpha P+\bm{\omega}A)t_{1}={\bm{r}}_{1}$ to find $t_{1}.$
$2:$ Solve $(\beta Q+\bm{\omega}D)t_{2}={\bm{r}}_{2}$ to find $t_{2}.$
$3:$ Solve $\widehat{R}{\bm{z}}_{3}={\bm{r}}_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain ${\bm{z}}_{3}.$
$4:$ Solve $(\alpha P+\bm{\omega}A){\bm{z}}_{1}={\bm{r}}_{1}-B^{T}{\bm{z}}_{3}$ to find ${\bm{z}}_{1}.$
$5:$ Solve $(\beta Q+\bm{\omega}D){\bm{z}}_{2}={\bm{r}}_{2}-\bm{\omega}C{\bm{z}}_{3}$ to obtain ${\bm{z}}_{2}.$
Remark 3.1.
Algorithm 1 necessitates solving two linear subsystems having the coefficient matrix $(\alpha P+\omega A)$, two subsystems having the coefficient matrix $(\beta Q+\bm{\omega}D),$ and one subsystem with the coefficient matrix $\widehat{R}.$ Since the coefficient matrices are SPD matrices, we have the flexibility to employ exact solvers, such as Cholesky factorization, and inexact solvers, like the preconditioned conjugate gradient method, to solve them efficiently. Nevertheless, to solve steps $1$ and $4$, only one Cholesky factorization of $\alpha P+\bm{\omega}A$ and to solve steps $2$ and $5$, only one Cholesky factorization of $\beta Q+\bm{\omega}D$ is needed to perform.
4 Convergence analysis of the GSS iterative method
The purpose of this section is to investigate the convergence behavior of the proposed GSS iterative method. To achieve this, the following result plays a crucial role.
Lemma 4.1.
Let $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row matrices. Then the matrix $\Theta^{-1}\mathcal{B}$ is positive stable.
Proof.
Since $\Theta$ is SPD, $\Theta^{-1}\mathcal{B}$ is similar to the matrix $\Theta^{-\frac{1}{2}}\mathcal{B}\Theta^{-\frac{1}{2}}.$ By computation, we find that the block structure of the matrix $\Theta^{-\frac{1}{2}}\mathcal{B}\Theta^{-\frac{1}{2}}$ is identical to that of $\mathcal{B}$. Therefore, using Proposition 2.3, it follows that $\Theta^{-\frac{1}{2}}\mathcal{B}\Theta^{-\frac{1}{2}},$ and hence, $\Theta^{-1}\mathcal{B}$ is positive stable. $\blacksquare$
As noted in [35], a stationary iterative method in the form (3.4) converges if and only if the iteration matrix has the spectral radius strictly less than one. The following result discusses the convergence of the GSS iterative method (3.4).
Theorem 4.1.
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row matrices. Let $\alpha,\beta,\tau,\bm{\omega}>0,$ $P\in{\mathbb{R}}^{n\times n},Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}$ be SPD matrices. Then, the GSS iterative method converges to the unique solution of the DSPP (1.1) if
${\displaystyle{\bm{\omega}\geq\max\left\{\frac{1}{2}-\frac{\lambda_{\min}(\widetilde{\Theta})}{\vartheta(\Theta^{-1}\mathcal{B})^{2}},0\right\}}},$
where $\widetilde{\Theta}={\displaystyle{\frac{\Theta^{-1}\mathcal{B}+\mathcal{B}^{T}\Theta^{-1}}{2}}}.$
Proof.
The iteration matrix of the GSS iterative method (3.4) is
$\mathcal{G}=\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{N}_{\mathrm{GSS}}$ $=(\Theta+\bm{\omega}\mathcal{B})^{-1}(\Theta-(1-\bm{\omega})\mathcal{B})$ (4.1)
$=(I-\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}).$ (4.2)
Let $\lambda$ be an eigenvalue of $\mathcal{G}$. Then
$(I-\bm{\omega}\Theta^{-1}\mathcal{B})^{-1}\left(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}\right){\bf x}=\lambda{\bf x},$ (4.3)
where ${\bf x}\in{\mathbb{R}}^{n+m+l}$ is the corresponding eigenvector. From (4.3), we can write
$(I-(1-\bm{\omega})\Theta^{-1}\mathcal{B}){\bf x}=\lambda(I-\bm{\omega}\Theta^{-1}\mathcal{B}){\bf x}$ (4.4)
$\implies$ $((1-\bm{\omega})+\lambda\bm{\omega})\Theta^{-1}\mathcal{B}{\bf x}=(1-\lambda){\bf x}.$ (4.5)
Note that $\lambda\neq 1,$ otherwise (4.5) reduces to $\Theta^{-1}\mathcal{B}{\bf x}=0,$ which implies that ${\bf x}=0.$ On the other hand, $(1-\bm{\omega})+\lambda\bm{\omega}\neq 0,$ otherwise $(1-\lambda){\bf x}=0.$ This gives ${\bf x}=0,$ which contradicts the assumption that ${\bf x}$ is an eigenvector. Therefore, from (4.5) we get
$\Theta^{-1}\mathcal{B}{\bf x}=\frac{1-\lambda}{(1-\bm{\omega})+\lambda\bm{\omega}}{\bf x}.$ (4.6)
Thus $\theta:=\frac{1-\lambda}{(1-\bm{\omega})+\lambda\bm{\omega}}$ is an eigenvalue of $\Theta^{-1}\mathcal{B}.$ Further, we can write
$\lambda=\frac{1-(1-\bm{\omega})\theta}{1+\bm{\omega}\theta}.$
Therefore, $|\lambda|<1$ if and only if $|1-(1-\bm{\omega})\theta|<|1+\bm{\omega}\theta|,$ i.e.,
$(1-(1-\bm{\omega})\mathfrak{R}(\theta))^{2}+(1-\bm{\omega})^{2}\Im(\theta)^{2}<(1+\bm{\omega}\mathfrak{R}(\theta))^{2}+\bm{\omega}^{2}\Im(\theta).$ (4.7)
Consequently, it follows from (4.7) that the iterative method (3.4) is convergent if
$02\mathfrak{R}(\theta)+(2\bm{\omega}-1)|\theta|^{2}>$ 0 . (4.8)
By Lemma 4.1, we have $\mathfrak{R}(\theta)>0,$ and this implies $\frac{\mathfrak{R}(\theta)}{|\theta|^{2}}>0.$ From (4.8), we get $\bm{\omega}>\frac{1}{2}-\frac{\mathfrak{R}(\theta)}{|\theta|^{2}}.$
Next, assume that $\bm{w}$ is the eigenvector corresponding to the eigenvalue $\theta.$ Then $\Theta^{-1}\mathcal{B}\bm{w}=\theta\bm{w}.$ Multiplying by $\bm{w}^{*}$ from the left side, we have $\bm{w}^{*}\Theta^{-1}\mathcal{B}\bm{w}=\theta\bm{w}^{*}\bm{w},$ and taking conjugate transpose gives $\bm{w}^{*}\mathcal{B}^{T}\Theta^{-1}\bm{w}=\bar{\theta}\bm{w}^{*}\bm{w}.$ Hence,
$\mathfrak{R}(\theta)=\frac{\bm{w}^{*}(\Theta^{-1}\mathcal{B}+\mathcal{B}^{T}\Theta^{-1})\bm{w}}{2\bm{w}^{*}\bm{w}}\geq\lambda_{\min}(\widetilde{\Theta}).$
Again $|\theta|\leq\vartheta(\Theta)$ gives ${\displaystyle{\frac{1}{2}-\frac{\mathfrak{R}(\theta)}{|\theta|^{2}}\leq\frac{1}{2}-\frac{\lambda_{\min}(\widetilde{\Theta})}{\vartheta(\Theta)^{2}}}}.$ Since $\bm{\omega}>0,$ the GSS iterative method is convergent if ${\displaystyle{\bm{\omega}>\max\left\{\frac{1}{2}-\frac{\lambda_{\min}(\widetilde{\Theta})}{\vartheta(\Theta)^{2}},0\right\}}}.$blacksquare
Remark 4.1.
Note that if $\bm{\omega}\geq\frac{1}{2},$ then the condition (4.7) holds. This shows that the GSS iterative method (3.4) is convergent for any initial guess vector if $\bm{\omega}\geq\frac{1}{2}.$
Notably, solving the DSPP (1.1) is same as to solve the preconditioned linear system $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}{\bf u}=\mathcal{M}_{\text{GSS}}^{-1}{\bm{b}}$. Hence, as an immediate consequence of Theorem 4.1 and Remark 4.1, we have the following results regarding the clustering properties of the spectrum of the preconditioned matrix $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}$.
Theorem 4.2.
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{p\times p}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{p\times m}$ are full row matrices. Let $\mathcal{M}_{\mathrm{GSS}}$ be defined as in (3.2) and $\lambda$ be an eigenvalue of the preconditioned matrix $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$. If $\bm{\omega}\geq 1/2,$ then $\lambda$ satisfies the following:
$|\lambda-1|<1,$
i.e., all eigenvalues of the preconditioned matrix $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$ are entirely contained in a circle centered at $(1,0)$ with radius strictly less one.
Proof.
The proof follows immediately from the identity $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B}=\mathcal{M}_{\text{GSS}}^{-1}(\mathcal{M}_{\text{GSS}}-\mathcal{N}_{\text{GSS}})=I-\mathcal{G}.$blacksquare
5 Two relaxed variants of GSS preconditioner
To enhance efficiency, this section introduces two relaxed variants of the GSS preconditioner. By removing $\alpha P$ from $(1,1)$ block and $\beta Q$ from $(2,2)$ block of $\mathcal{M}_{\mathrm{GSS}},$ we obtain the following two relaxed GSS (RGSS) preconditioners:
$\mathcal{M}_{\text{RGSS-I}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right]$ (5.1)
and
$\mathcal{M}_{\text{RGSS-II}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{\omega}B^{T}\\ \bm{0}&\bm{\omega}D&\bm{\omega}C\\ -\bm{\omega}B&-\bm{\omega}C^{T}&\tau R\end{matrix}\right].$ (5.2)
In the implementation, at each step of RGSS preconditioners in conjunction with GMRES, we are required to solve the linear systems of the following forms:
$\mathcal{M}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}\leavevmode\nobreak\ \leavevmode\nobreak\ \text{or}\leavevmode\nobreak\ \leavevmode\nobreak\ \mathcal{M}_{\text{RGSS-II}}{\bm{w}}={\bm{r}}.$ (5.3)
Set ${\mathcal{R}}_{1}=\tau R+\bm{\omega}BA^{-1}B^{T}+\bm{\omega}^{2}C^{T}(\beta Q+\bm{\omega}D)^{-1}C$ and $\mathcal{R}_{2}=\tau R+\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C,$ then
$\mathcal{M}_{\text{RGSS-I}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-\bm{\omega}C^{T}(\beta Q+\bm{\omega}D)^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\beta Q+\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{1}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{\omega}(\beta Q+\bm{\omega}D)^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.4)
$\mathcal{M}_{\text{RGSS-II}}=$ $\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&-C^{T}D^{-1}&I\end{matrix}\right]\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D&\bm{0}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&D^{-1}C\\ \bm{0}&\bm{0}&I\end{matrix}\right],$ (5.5)
By applying a similar methodology as in Algorithm 1, we derive Algorithms 2 and 3 for implementing the RGSS preconditioners.
Algorithm 2 Solving $\mathcal{M}_{\text{RGSS-I}}{\bm{w}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\beta,\tau,\bm{\omega},$ SPD matrices $Q\in{\mathbb{R}}^{l\times l}$ and $R\in{\mathbb{R}}^{m\times m}.$
Output: Solution vector ${\bm{w}}=[\bm{w}_{1}^{T},\bm{w}_{2}^{T},\bm{w}_{3}^{T}]^{T}\in{\mathbb{R}}^{n+l+m}.$
Steps:
$1:$ Solve $At_{1}=r_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $(\beta Q+\bm{\omega}D)t_{2}=r_{2}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{1}\bm{w}_{3}=r_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(r_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $(\beta Q+\bm{\omega}D)\bm{w}_{2}=r_{2}-\bm{\omega}C\bm{w}_{3}$ to obtain $\bm{w}_{2}.$
Algorithm 3 Solving $\mathcal{M}_{\text{RGSS-II}}{\bm{w}}={\bm{r}}$
Input: The matrices $A,B,C,D,$ $\bm{r},$ positive parameters $\tau,\bm{\omega}$ and the SPD matrix $R\in{\mathbb{R}}^{m\times m}.$
Output: Solution vector ${\bm{w}}=[\bm{w}_{1}^{T},\bm{w}_{2}^{T},\bm{w}_{3}^{T}]^{T}\in{\mathbb{R}}^{n+l+m}.$
Steps:
$1:$ Solve $At_{1}=r_{1}/\bm{\omega}$ to find $t_{1}.$
$2:$ Solve $Dt_{2}=r_{2}/\bm{\omega}$ to find $t_{2}.$
$3:$ Solve $\mathcal{R}_{2}\bm{w}_{3}=r_{3}+\bm{\omega}Bt_{1}+\bm{\omega}C^{T}t_{2}$ to obtain $\bm{w}_{3}.$
$4:$ Solve $A\bm{w}_{1}=\frac{1}{\bm{\omega}}(r_{1}-B^{T}\bm{w}_{3})$ to find $\bm{w}_{1}.$
$5:$ Solve $D\bm{w}_{2}=\frac{1}{\bm{\omega}}(r_{2}-\bm{\omega}C\bm{w}_{3})$ to obtain $\bm{w}_{2}.$
Next, we investigate the spectral properties of the preconditioned matrices $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}.$
Theorem 5.1.
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $Q$ and $R$ be SPD matrices. Then, the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}{\text{-}}\mathrm{I}}^{-1}\mathcal{B}$ has ${\displaystyle{\frac{1}{\bm{\omega}}}}$ as the eigenvalue with multiplicity $n.$ Further, let $\mu$ be an eigenvalue among the remaining $m+l$ eigenvalues with the corresponding eigenvector $[u^{T},v^{T}]^{T}$ such that $\|\sqrt{\beta}Q^{1/2}u^{T},\sqrt{\tau}R^{1/2}v^{T}]^{T}\|_{2}=1$. Then
  1. (1)
    if $\Im(\frac{\mu}{1-\bm{\omega}\mu})\neq 0,$ we have $\beta\|Q^{1/2}u\|^{2}_{2}=\frac{1}{2}=\tau\|R^{1/2}v\|^{2}_{2}$ and
    $\mathfrak{R}\left(\frac{\mu}{1-\bm{\omega}\mu}\right)=\frac{1}{2}\left(\frac{u^{*}Du}{\beta u^{*}Qu}+\frac{v^{*}Sv}{\tau v^{*}Rv}\right).$
    Thus, it satisfies the following bounds:
    $\left\{\begin{array}[]{c}\frac{1}{2}\left(\frac{\lambda_{\min}(Q^{-1}D)}{\beta}+\frac{\lambda_{\min}(R^{-1}S)}{\tau}\right)\leq\mathfrak{R}\left(\frac{\mu}{1-\bm{\omega}\mu}\right)\leq\frac{1}{2}\left(\frac{\lambda_{\max}(Q^{-1}D)}{\beta}+\frac{\lambda_{\max}(R^{-1}S)}{\tau}\right),\\ |\Im\left(\frac{\mu}{1-\bm{\omega}\mu}\right)|\leq\sigma_{\max}\left(\frac{1}{\sqrt{\beta\tau}}R^{-\frac{1}{2}}CQ^{-\frac{1}{2}}\right).\end{array}\right.$ (5.8)
  2. (2)
    If $\Im(\frac{\mu}{1-\bm{\omega}\mu})=0,$ we have
    $\frac{\mu}{1-\bm{\omega}\mu}=\frac{u^{*}Du+v^{*}Sv}{\beta u^{*}Qu+\tau v^{*}Rv}$
    and it holds that:
    $\begin{array}[]{c}2\min\left\{\frac{\lambda_{\min}(Q^{-1}D)}{\beta},\frac{\lambda_{\min}(R^{-1}S)}{\tau}\right\}\leq\frac{\mu}{1-\bm{\omega}\mu}\leq\max\left\{\frac{\lambda_{\max}(Q^{-1}D)}{\beta},\frac{\lambda_{\max}(R^{-1}S)}{\tau}\right\}.\end{array}$ (5.10)
Proof.
Let $S=BA^{-1}B^{T}.$ Then the saddle point matrix $\mathcal{B}$ and the preconditioner $\mathcal{M}_{\text{RGSS-II}}$ admits the following decompositions:
$\mathcal{B}=\mathbf{L}\mathbf{D}\mathbf{U}\quad\text{and}\quad\mathcal{M}_{\text{RGSS-I}}=\mathbf{L}\widetilde{\mathbf{D}}\mathbf{U},$ (5.11)
where $\mathbf{L}=\left[\begin{matrix}I&\bm{0}&\bm{0}\\ \bm{0}&I&\bm{0}\\ -BA^{-1}&\bm{0}&I\end{matrix}\right],$ $\mathbf{D}=\left[\begin{matrix}A&\bm{0}&\bm{0}\\ \bm{0}&D&C\\ \bm{0}&-C^{T}&S\end{matrix}\right],$ $\widetilde{\mathbf{D}}=\left[\begin{matrix}\bm{\omega}A&\bm{0}&\bm{0}\\ \bm{0}&\bm{\omega}D+\beta Q&\bm{\omega}C\\ \bm{0}&-\bm{\omega}C^{T}&S+\tau R\end{matrix}\right]$ and $\mathbf{U}=\left[\begin{matrix}I&\bm{0}&A^{-1}B^{T}\\ \bm{0}&I&\bm{0}\\ \bm{0}&\bm{0}&I\end{matrix}\right].$ Using decompositions in (5.11), we obtain
$\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}=\mathbf{U}^{-1}\widetilde{\mathbf{D}}^{-1}\mathbf{D}\mathbf{U}.$ (5.12)
Therefore, $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ is similar to $\widetilde{\mathbf{D}}^{-1}\mathbf{D},$ which is given by
$\widetilde{\mathbf{D}}^{-1}\mathbf{D}=\left[\begin{matrix}\bm{\omega}^{-1}I&\bm{0}\\ \bm{0}&\mathbf{M}^{-1}{\mathbf{A}}\end{matrix}\right],$ (5.13)
where $\mathbf{M}=\left[\begin{matrix}\bm{\omega}D+\beta Q&\bm{\omega}C\\ -\bm{\omega}C^{T}&\bm{\omega}S+\tau R\end{matrix}\right]$ and $\mathbf{A}=\left[\begin{matrix}D&C\\ -C^{T}&S\end{matrix}\right].$ Hence, $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ has an eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity at least $n$ and while the remaining eigenvalues are those of the preconditioned matrix $\mathbf{M}^{-1}\mathbf{A}.$ Consider $\mathbf{I}=\left[\begin{matrix}\bm{0}&I\\ -I&\bm{0}\end{matrix}\right]\in{\mathbb{R}}^{l+m},$ then $\mathbf{I}\left(\mathbf{I}^{-1}\mathbf{M}\mathbf{I}\right)^{-1}\left(\mathbf{I}^{-1}\mathbf{A}\mathbf{I}\right)\mathbf{I}^{-1}.$ Consequently, $\mathbf{M}^{-1}\mathbf{A}$ is similar to $\widetilde{\mathbf{M}}^{-1}\widetilde{\mathbf{A}},$ where
$\widetilde{\mathbf{M}}:=\mathbf{I}^{-1}\mathbf{M}\mathbf{I}=\left[\begin{matrix}\bm{\omega}D+\beta Q&\bm{\omega}C^{T}\\ -\bm{\omega}C&\bm{\omega}S+\tau R\end{matrix}\right]\quad\text{and}\quad\widetilde{\mathbf{A}}:=\mathbf{I}^{-1}\mathbf{A}\mathbf{I}=\left[\begin{matrix}D&C^{T}\\ -C&S\end{matrix}\right].$ (5.14)
Let $\mu$ be an eigenvalue of the matrix $\widetilde{\mathbf{M}}^{-1}\widetilde{\mathbf{A}}$ with the corresponding eigenvalue $[u^{T},v^{T}]^{T}.$ Assert that $\mu=\frac{1}{\bm{\omega}},$ then
$\left[\begin{matrix}\beta Q&\bm{0}\\ \bm{0}&\tau R\end{matrix}\right]\left[\begin{matrix}u\\ v\end{matrix}\right]=\bm{0},$ (5.15)
which gives $[u^{T},v^{T}]^{T}=\bm{0}.$ This contradicts to the assumption that $[u^{T},v^{T}]^{T}$ is an eigenvector, and hence $\mu\neq\frac{1}{\bm{\omega}}.$ Since, $\mu$ is an eigenvalue of $\widetilde{\mathbf{M}}^{-1}\widetilde{\mathbf{A}}$, we have
$\mu\left[\begin{matrix}\bm{\omega}D+\beta Q&\bm{\omega}C^{T}\\ -\bm{\omega}C&\bm{\omega}S+\tau R\end{matrix}\right]\left[\begin{matrix}u\\ v\end{matrix}\right]=\left[\begin{matrix}D&C^{T}\\ -C&S\end{matrix}\right]\left[\begin{matrix}u\\ v\end{matrix}\right]$
$\Longrightarrow\frac{\mu}{1-\bm{\omega}\mu}\left[\begin{matrix}\beta Q&\bm{0}\\ \bm{0}&\tau R\end{matrix}\right]\left[\begin{matrix}u\\ v\end{matrix}\right]=\left[\begin{matrix}D&C^{T}\\ -C&S\end{matrix}\right]\left[\begin{matrix}u\\ v\end{matrix}\right]$
$\Longrightarrow\frac{\mu}{1-\bm{\omega}\mu}\left[\begin{matrix}\sqrt{\beta}Q^{\frac{1}{2}}u\\ \sqrt{\tau}R^{\frac{1}{2}}v\end{matrix}\right]=\left[\begin{matrix}\frac{1}{\sqrt{\beta}}Q^{-\frac{1}{2}}&\bm{0}\\ \bm{0}&\frac{1}{\sqrt{\beta}}R^{-\frac{1}{2}}\end{matrix}\right]\left[\begin{matrix}D&C^{T}\\ -C&S\end{matrix}\right]\left[\begin{matrix}\frac{1}{\sqrt{\beta}}Q^{-\frac{1}{2}}&\bm{0}\\ \bm{0}&\frac{1}{\sqrt{\tau}}R^{-\frac{1}{2}}\end{matrix}\right]\left[\begin{matrix}\sqrt{\beta}Q^{\frac{1}{2}}u\\ \sqrt{\tau}R^{\frac{1}{2}}v\end{matrix}\right]$
$\Longrightarrow\frac{\mu}{1-\bm{\omega}\mu}\left[\begin{matrix}\widetilde{u}\\ \widetilde{v}\end{matrix}\right]=\left[\begin{matrix}\widetilde{D}&\widetilde{C}^{T}\\ -\widetilde{C}&\widetilde{S}\end{matrix}\right]\left[\begin{matrix}\widetilde{u}\\ \widetilde{v}\end{matrix}\right],$
where $\widetilde{D}=\frac{1}{\beta}Q^{-\frac{1}{2}}DQ^{-\frac{1}{2}},$ $\widetilde{C}=\frac{1}{\sqrt{\beta\tau}}R^{-\frac{1}{2}}CQ^{-\frac{1}{2}},$ $\widetilde{S}=\frac{1}{\tau}R^{-\frac{1}{2}}SR^{-\frac{1}{2}},$ $\widetilde{u}=\sqrt{\beta}Q^{\frac{1}{2}}u$ and $\widetilde{v}=\sqrt{\tau}R^{\frac{1}{2}}v.$ Since, $\widetilde{D}$ and $\widetilde{S}$ are SPD matrices and $\widetilde{C}$ has full row rank, according to Proposition $2.12$ in [9], we have the desired bounds of (5.8) and (5.10). $\blacksquare$
Remark 5.1.
The asymptotic behavior of the eigenvalue $\mu$ is analyzed according to Theorem 5.1 when the iteration parameters $\beta$ and $\tau$ approach zero from the positive side. This analysis is conducted in the following two cases:
  • Let $\theta=\frac{\mu}{1-\bm{\omega}\mu}.$ When $\Im(\theta)\neq 0,$ we have $\mathfrak{R}(\theta)\rightarrow+\infty$ and $|\Im(\theta)|\rightarrow+\infty$ as $\beta,\tau\rightarrow 0_{+}.$ Then, it holds that
    $\mu=\left[\frac{1}{\bm{\omega}}-\frac{1+\bm{\omega}\mathfrak{R}(\theta)}{(1+\bm{\omega}\mathfrak{R}(\theta))^{2}+\bm{\omega}^{2}\Im(\theta)^{2}}\right]+{i\mkern 1.0mu}\frac{\Im(\theta)}{(1+\bm{\omega}\mathfrak{R}(\theta))^{2}+\bm{\omega}^{2}\Im(\theta)^{2}}\rightarrow\frac{1}{\bm{\omega}}\leavevmode\nobreak\ \text{as}\leavevmode\nobreak\ \beta,\tau\rightarrow 0_{+}.$
  • When $\Im(\theta)=0,$ we have $\theta\rightarrow+\infty$ as $\beta,\tau\rightarrow 0_{+}.$ Then, it holds that
    $\mu=\frac{\theta}{1+\bm{\omega}\theta}\rightarrow\frac{1}{\bm{\omega}},\leavevmode\nobreak\ \text{as}\leavevmode\nobreak\ \beta,\tau\rightarrow 0_{+}.$ (5.16)
The following theorem establishes the spectral properties for the RGSS-II preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$.
Theorem 5.2.
Assume that $A$ and $D$ are SPD matrices, $B$ and $C$ have full row rank, and let $Q$ be an SPD matrix. Then the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\frac{1}{\bm{\omega}}}}$ with multiplicity $n+l.$ The remaining eigenvalues satisfies the generalized eigenvalue problem $(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}=\lambda\mathcal{R}_{2}{\bf x},$ where $\mathcal{R}_{2}=\tau R+\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C.$
Proof.
From (5.5), we have
$\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ $=\left[\begin{matrix}\bm{\omega}^{-1}A&\bm{0}&-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{\omega}^{-1}D&-D^{-1}C\mathcal{R}_{2}^{-1}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}\end{matrix}\right]\left[\begin{matrix}A&\bm{0}&B^{T}\\ \bm{0}&D&C\\ \bm{0}&\bm{0}&\mathbf{X}\end{matrix}\right]$
$=\left[\begin{matrix}\bm{\omega}^{-1}I&\bm{0}&\bm{\omega}^{-1}AB^{T}-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\mathbf{X}\\ \bm{0}&\bm{\omega}^{-1}I&\bm{\omega}^{-1}DC-D^{-1}C\mathcal{R}_{2}^{-1}\mathbf{X}\\ \bm{0}&\bm{0}&\mathcal{R}_{2}^{-1}\mathbf{X}\end{matrix}\right],$ (5.17)
where $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R.$ Thus, $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ has the eigenvalue ${\displaystyle{\lambda=\frac{1}{w}}}$ with multiplicity at least $n+l,$ and the rest of eigenvalues satisfy the generalized eigenvalue problem
$(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}=\lambda\mathcal{R}_{2}{\bf x}.$
Hence, the proof is completed. $\blacksquare$
Corollary 5.1.
Suppose that assumption on Theorem 5.2 holds. Then the eigenvalues of $\mathcal{M}_{\mathrm{RGSS}\text{-}\mathrm{II}}^{-1}\mathcal{B}$ satisfies
$\lambda\in[\Lambda_{\min},\Lambda_{\max}],$ (5.18)
where
$\Lambda_{\min}=\frac{\eta_{\min}+\xi_{\min}+\tau}{\bm{\omega}\eta_{\max}+\bm{\omega}\xi_{\max}+\tau},\leavevmode\nobreak\ \Lambda_{\max}=\frac{\eta_{\max}+\xi_{\max}+\tau}{\bm{\omega}\eta_{\min}+\bm{\omega}\xi_{\min}+\tau},$
$\eta_{\min}=\lambda_{\min}(R^{-1}BA^{-1}B^{T}),$ $\eta_{\max}=\lambda_{\max}(R^{-1}BA^{-1}B^{T}),$ $\xi_{\min}=\lambda_{\min}(R^{-1}C^{T}D^{-1}C)$ and $\xi_{\max}=\lambda_{\max}(R^{-1}C^{T}D^{-1}C).$
Proof.
Premultiplying by ${\bf x}^{T}$ of the generalized eigenvalue problem $(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}=\lambda\mathcal{R}_{2}{\bf x},$ we obtain
$\lambda$ $=\frac{{\bf x}^{T}(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau R){\bf x}}{{\bf x}^{T}(\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C+\tau R){\bf x}}$
$=\frac{{(R^{1/2}\bf x)}^{T}R^{-1/2}(BA^{-1}B^{T}+C^{T}D^{-1}C+\tau)R^{-1/2}{(R^{1/2}\bf x)}}{{(R^{1/2}\bf x)}^{T}R^{-1/2}(\bm{\omega}BA^{-1}B^{T}+\bm{\omega}C^{T}D^{-1}C+\tau)R^{-1/2}{(R^{1/2}\bf x)}}.$ (5.19)
Since $R^{-1/2}BA^{-1}B^{T}R^{-1/2}$ is similar to $R^{-1}BA^{-1}B^{T}$ and $R^{-1/2}C^{T}A^{-1}CR^{-1/2}$ is similar to $R^{-1}C^{T}A^{-1}C,$ and $\lambda_{\min}(X)\leq\theta\leq\lambda_{\max}(X)$ for any $\theta\in\sigma(X),$ where $X$ is SPD, the proof follows from (5). $\blacksquare$
Next, we will examine the properties of the minimal polynomial of the preconditioned matrix $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B},$ which is crucial for effective Krylov subspace acceleration.
Theorem 5.3.
Assume that $A\in{\mathbb{R}}^{n\times n}$ and $D\in{\mathbb{R}}^{l\times l}$ are SPD matrices, $B\in{\mathbb{R}}^{m\times n}$ and $C\in{\mathbb{R}}^{l\times m}$ are full row rank matrices. Then, the degree of the minimal polynomial of the preconditioned matrix $\mathcal{M}_{\mathrm{RGSS-II}}^{-1}\mathcal{B}$ is almost $m+1.$ Therefore, the dimension of the Krylov subspace $\mathcal{K}(\mathcal{M}_{\mathrm{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ is almost $m+1.$
Proof.
From (5), we obtain
$\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}=\left[\begin{matrix}w^{-1}I&\bm{0}&\Sigma_{1}\\ \bm{0}&w^{-1}I&\Sigma_{2}\\ \bm{0}&\bm{0}&\Sigma_{3}\end{matrix}\right],$ (5.20)
where $\Sigma_{1}=\bm{\omega}^{-1}AB^{T}-A^{-1}B^{T}\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{2}=\bm{\omega}^{-1}DC-D^{-1}C\mathcal{R}_{2}^{-1}\mathbf{X},$ $\Sigma_{3}=\mathcal{R}_{2}^{-1}\mathbf{X},$ and $\mathbf{X}=BA^{-1}B^{T}+C^{T}D^{-1}C+R.$ Let $\mu_{i},$ $i=1,2,\ldots,m$ be the eigenvalues of of $\Sigma_{3}.$ Then they are also eigenvalues of the preconditioned matrix $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}.$ Then the characteristic polynomial of $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ is given by
$f(\lambda)=\left(\lambda-\frac{1}{w}\right)^{n+l}\prod_{i=1}^{m}(\lambda-\mu_{i}).$
Consider the polynomial ${g(\lambda)=(\lambda-\frac{1}{\bm{\omega}})\prod\limits_{i=1}^{m}(\lambda-\mu_{i})}.$ Then
$g(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B})$ $=(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}-\frac{1}{\bm{\omega}}I)\prod\limits_{i=1}^{m}(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}-\mu_{i}I)$
$=\left[\begin{matrix}\bm{0}&\bm{0}&\Sigma_{1}\\ \bm{0}&\bm{0}&\Sigma_{2}\\ \bm{0}&\bm{0}&\Sigma_{3}-\bm{\omega}^{-1}I\end{matrix}\right]\left[\begin{matrix}{\prod\limits_{i=1}^{n}(\bm{\omega}^{-1}I-\mu_{i}I)}&\bm{0}&\Sigma_{1}\\ \bm{0}&{\prod\limits_{i=1}^{l}(\bm{\omega}^{-1}I-\mu_{i}I)}&\Sigma_{2}\\ \bm{0}&\bm{0}&{\prod\limits_{i=1}^{m}(\Sigma_{3}-\mu_{i}I)}\end{matrix}\right].$ (5.21)
Given that $\Sigma_{3}$ has the eigenvalues $\mu_{i},$ $i=1,2,\ldots,m,$ we obtain ${\prod\limits_{i=1}^{m}(\Sigma_{3}-\mu_{i}I)}=\bm{0}.$ Therefore, from (5.21), we obtain $g(\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B})=\bm{0}.$ Hence, by Cayley-Hamilton theorem, the degree of the minimal polynomial of $\mathcal{M}_{\text{GSS-II}}^{-1}\mathcal{B}$ is almost $m+1.$
As mentioned in [35], the degree of the minimal polynomial of a matrix and the dimension of the associated Krylov subspace are equal. Therefore, the Krylov subspace $\mathcal{K}(\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B},{\bm{b}})$ has dimension atmost $m+1.$ Hence, the proof follows. $\blacksquare$
Based on the property outlined in Theorem 5.3, Krylov subspace methods like PGMRES require maximum $m+1$ iterations to solve the system (1.1).
6 Discussion on the selection of the parameters
It is worth noting that proposed GSS, RGSS-I and RGSS-II preconditioners involve the parameters $\alpha,$ $\beta,$ $\tau$ and $\bm{\omega}.$ As a general criterion for a preconditioner to perform efficiently, it should be as close as possible to the coefficient matrix of the system [10], we find the optimal choices for the proposed GSS preconditioner by minimizing $\|\mathcal{N}_{\text{GSS}}\|_{F}=\|\mathcal{M}_{\text{GSS}}-\mathcal{B}\|_{F}.$ First, we define the function $\varphi$ by
$\varphi(\alpha,\beta,\tau,\bm{\omega})=\|\mathcal{N}_{\text{GSS}}\|_{F}^{2}=tr(\mathcal{N}_{\text{GSS}}^{T}\mathcal{N}_{\text{GSS}})>0.$
After some easy calculations, we obtain
$\varphi(\alpha,\beta,\tau,\bm{\omega})=$ $\alpha^{2}\|P\|_{F}^{2}+\beta^{2}\|Q\|_{F}^{2}+\tau^{2}\|R\|_{F}^{2}+(1-\bm{\omega})^{2}\|A\|_{F}^{2}+(1-\bm{\omega})^{2}\|D\|_{F}^{2}$
$+2\alpha(\bm{\omega}-1)tr(PA)+2\beta(\bm{\omega}-1)tr(QD)+2(1-\bm{\omega})^{2}\|B\|_{F}^{2}$
$+2(1-\bm{\omega})^{2}\|C\|_{F}^{2}.$
Now we need to select the parameters $\alpha,\beta,\tau$ and $\bm{\omega}$ such that $\varphi(\alpha,\beta,\tau,\bm{\omega})$ is very small. Since
$\lim_{\alpha,\,\beta,\,\tau,\,\bm{\omega}\rightarrow 0_{+}}\varphi(\alpha,\beta,\tau,\bm{\omega})=(1-\bm{\omega})^{2}\|A\|_{F}^{2}+(1-\bm{\omega})^{2}\|D\|_{F}^{2}+2(1-\bm{\omega})^{2}\|B\|_{F}^{2}+2(1-\bm{\omega})^{2}\|C\|_{F}^{2},$
we can select $\bm{\omega}=1$ and $\alpha,\,\beta,\,\tau,\,\bm{\omega}\rightarrow 0_{+}$ such that $\varphi(\alpha,\beta,\tau,\bm{\omega})\rightarrow 0_{+},$ and consequently, $\mathcal{N}_{\text{GSS}}\rightarrow\bm{0}.$
In the sequel, the GSS preconditioner is equivalently rewritten as: $\mathcal{M}_{\text{GSS}}=\bm{\omega}\widetilde{\mathcal{M}}_{\text{GSS}},$ where
~$\widetilde{\mathcal{M}}_{\text{GSS}}=\left[\begin{matrix}\frac{\alpha}{\bm{\omega}}P+A&\bm{0}&B^{T}\\ \bm{0}&\frac{\beta}{\bm{\omega}}Q+D&C\\ -B&-C^{T}&\frac{\tau}{\bm{\omega}}R\end{matrix}\right]$
can be regarded as a scaled preconditioner. A preconditioner is considered efficient if it closely approximates the coefficient matrix, and for given $\alpha,\beta,\tau>0,$$\widetilde{\mathcal{M}}_{\text{GSS}}-\mathcal{B}=\frac{1}{\bm{\omega}}\Theta\rightarrow\bm{0}$ as $\bm{\omega}\rightarrow\infty.$ Similar studies apply to the RGSS-I and RGSS-II preconditioners as well. The performance of the proposed preconditioners by varying the parameters is shown in the numerical experiment section.
7 Numerical experiments
To demonstrate the effectiveness and robustness of the proposed preconditioners GSS, RGSS-I and RGSS-II over the existing ones within the Krylov subspace iterative method to solve the DSPP, in this section, we perform a few numerical experiments. We compare our proposed GSS, RGSS-I and RGSS-II PGMRES methods (abbreviated as “GSS”, “RGSS-I” and “‘RGSS-II”, respectively) with the GMRES method and PGMRES methods with BD [13], DS [8], RDF [11], and SS [18] preconditioners (abbreviated as “BD”, “DS”, “RDF”, “SDF” and “SS”, respectively). The numerical results are presented in terms of iteration counts (abbreviated as “IT”) and elapsed CPU time in seconds (abbreviated as “CPU”).
The initial guess vector is set to ${\bm{u}}^{(0)}=\bm{0}\in{\mathbb{R}}^{n+l+m}$ for all iterative methods, and the method terminates if
${\tt RES}:=\frac{\|\mathcal{B}{\bm{u}}^{(k+1)}-{\bm{b}}\|_{2}}{\|{\bm{b}}\|_{2}}<10^{-6}$
or if the maximum number of iterations exceeds $5000$. All the linear subsystems involved in Algorithms 1, 2 and 3 are solved using the Cholesky Factorization. Numerical experiments are conducted in MATLAB R2024a on a Windows 11 system, using an Intel(R) Core(TM) i7-10700 CPU at 2.90 GHz with 16 GB of memory.
Example 1.
[13, 33] The Poisson control problem: Consider the distributed control problem (PDE-constrained optimization problem) defined by:
$\begin{array}[]{cc}&\min_{u,f}\frac{1}{2}\|u-\hat{u}\|^{2}_{L_{2}(\Omega)}+\frac{\bm{\beta}}{2}\|f\|^{2}_{L_{2}(\Omega)}\\ \text{such that }&-\nabla^{2}u=f\leavevmode\nobreak\ \text{in}\leavevmode\nobreak\ \Omega,\\ &\quad\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ u=g\leavevmode\nobreak\ \text{on}\leavevmode\nobreak\ \partial\Omega,\end{array}$ (7.4)
where $u$ is the state, and $\hat{u}$ is the desired state, $f$ is the control, $\Omega=[0,1]\times[0,1]$ is the domain with the boundary $\partial\Omega,$ and $0<{\bm{\beta}}\ll 1$ is the regularization parameter. By discretizing (7.4) using the Galerkin finite element method and then applying Lagrange multiplier techniques, we obtain the following system:
$\left[\begin{matrix}{\bm{\beta}}M&\bm{0}&K^{T}\\ \bm{0}&M&-M\\ -K&M&\bm{0}\end{matrix}\right]\left[\begin{matrix}u\\ f\\ \bm{\lambda}\end{matrix}\right]=\left[\begin{matrix}\bm{0}\\ b\\ -d\end{matrix}\right],$ (7.5)
where $M\in{\mathbb{R}}^{n\times n}$ and $K\in{\mathbb{R}}^{n\times n}$ are SPD mass matrix and discrete Laplacian, respectively. Note that by setting $A=\bm{\beta}M$, $B=K$, $C=-M$, and $\mathbf{b}=[\mathbf{0},b^{T},-d^{T}]^{T}$, (7.5) can be expressed in the form of the DSPP (1.1), where $n=m=p$.
The MATLAB code downloaded from [31], generates the linear system of the form (7.5) by using the following setup: parameters in “set_def_setup.m” are selected as def_setup.bc=‘dirichlet’, def_setup.beta=$1e-2$, def_setup.ob=1, def_setup.type=‘dist2d’ and def_setup.pow = $5$, $6$ and $7$.
Table 1: Experimental results of GMRES, BD, DS, RDF, SS, GSS, RGSS-I and RGSS-II PGMRES methods for Example 1 when ${\bm{\beta}}=0.1$
Process def__\__setup.pow $=5$ def_setup.pow $=6$ def_setup.pow $=7$
$\mathcal{B}$) $2883$ $11907$ $48378$
GMRES IT $579$ $2149$bf{--}- -
CPU $2.9618$ $367.9274$bf{--}- -
BD IT $10$ $10$ $10$
CPU $2.1418$ $40.6138$ $1512.0707$
DS IT $31$ $40$ $51$
CPU $2.8514$ $54.3105$ $1706.9543$
RDF IT $6$ $6$ $8$
CPU $0.8733$ $8.1585$ $302.9091$
SS IT $16$ $22$ $46$
CPU $2.12460$ $45.9988$ $3985.6141$
GSS IT $2$ $2$ $2$
$\bm{\omega}_{exp}=30$ CPU $0.6909$ $7.3233$ $226.8625$
RGSS-I IT $2$ $2$ $2$
$\bm{\omega}_{exp}=25$ CPU $0.7796$ $7.72871$ $249.7657$
RGSS-II IT $2$ $2$ $2$
$\bm{\omega}_{exp}=30$ CPU $0.7946$ $6.8447$ $238.5576$
[…]
Table 2: Experimental results of GMRES, BD, DS, RDF, SS, GSS, RGSS-I and RGSS-II PGMRES methods for Example 1 when ${\bm{\beta}}=0.001$
Process def_setup.pow $=5$ def_setup.pow $=6$ def_setup.pow $=7$
$\mathcal{B}$) $2883$ $11907$ $48378$
GMRES IT $919$ $2254$bf{--}- -
CPU $5.8702$ $409.5427$bf{--}- -
BD IT $19$ $19$ $19$
CPU $4.3498$ $84.9364$ $3034.4227$
DS IT $29$ $39$ $46$
CPU $2.6072$ $47.8576$ $1603.9128$
RDF IT $14$ $11$ $8$
CPU $1.4320$ $14.3846$ $324.8007$
SS IT $27$ $38$ $68$
CPU $3.3960$ $79.7055$ $5481.9674$
GSS IT $2$ $2$ $2$
$\bm{\omega}_{exp}=30$ CPU $0.7159$ $7.4041$ $242.8123$
RGSS-I IT $2$ $2$ $2$
$\bm{\omega}_{exp}=30$ CPU $0.78675$ $7.6515$ $221.0918$
RGSS-II IT $2$ $2$ $2$
$\bm{\omega}_{exp}=26$ CPU $0.6828$ $6.9229$ $247.4777$
[…]
(a) $\mathcal{B}$
(b) $\mathcal{P}_{\text{BD}}^{-1}\mathcal{B}$
(c) $\mathcal{P}_{\text{DS}}^{-1}\mathcal{B}$
(d) $\mathcal{P}_{\text{RDF}}^{-1}\mathcal{B}$
(e) $\mathcal{P}_{\text{SS}}^{-1}\mathcal{B}$
(f) $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B}$
(g) $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$
(h) $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$
Figure 1: Eigenvalue distributions of $\mathcal{B},$ $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\mathrm{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ for def__\__set.pow $=5$ with $\bm{\beta}=0.1$ for Example 1.
Parameter selection: For the DS preconditioner, we choose the parameter $\alpha$ (denoted by $\alpha_{\text{DS}}$) as follow [8]:
$\alpha_{\text{DS}}=\frac{\sqrt{tr(A^{T}A)+2tr(BB^{T})}+\sqrt{tr(D^{T}D)+2tr(C^{T}C)}}{2(n+m+l)}.$
For the RDF preconditioner, the parameter $\alpha$ is selected from the interval $(0,1)$ with a step size of 0.01. The optimal performance, in terms of minimal CPU times, is achieved with smaller values of $\alpha$ as found in [11]. For the SS preconditioner, we take $\alpha=0.01.$ For the GSS, RGSS-I and RGSS-II preconditioners parameters are selected as follows: $\alpha_{\text{DS}}=\beta=0.01,$ $\tau=0.001,$ $P=A,$ $Q=CC^{T}$ and $R=I.$ The optimal parameter $\bm{\omega}$ (denoted by $\bm{\omega}_{exp}$) is determined experimentally within the interval $[2,30]$ with step size one, which yields minimal CPU times.
Numerical results: The numerical results for GMRES and various PGMRES methods with $\bm{\beta}=0.1$ and $0.001$ are presented in Tables 1 and 2. We observe that the GMRES method exhibits a significantly slower convergence rate compared to all other PGMRES methods, even does not converge within $5000$ iterations when def_setup.pow $=7$. On the other hand, we observe that our proposed preconditioners outperform all the compared preconditioners in terms of both IT and CPU times. For the DS and SS preconditioners, the IT increases as the size of the saddle point matrix $\mathcal{B}$ grows. Whereas the proposed GSS, RGSS-I and RGSS-II preconditioners maintain a consistent IT regardless of matrix size.
Eigenvalue distributions: In order to better illustrate the superiority of the proposed GSS, RGSS-I and RGSS-II preconditioners, the eigenvalue distribution of $\mathcal{B}$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ are displayed in Figure 1. From Figure 1, we observe that the eigenvalues of the preconditioned matrices $\mathcal{M}_{\text{GSS}}^{-1}\mathcal{B},$ $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B},$ and $\mathcal{M}_{\text{RGSS-II}}^{-1}\mathcal{B}$ have clustered better than the coefficient matrix $\mathcal{B},$ and preconditioned matrices $\mathcal{P}_{\mathrm{BD}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{DS}}^{-1}\mathcal{B},$ $\mathcal{P}_{\mathrm{RDF}}^{-1}\mathcal{B}$ and $\mathcal{P}_{\mathrm{SS}}^{-1}\mathcal{B}.$ This indicates enhanced computational efficiency, highlighting the superiority of the GSS, RGSS-I, and RGSS-II preconditioners over existing methods.
Influence of the parameters ${\alpha,\beta,\tau,\bm{\omega}}$:
(a) By varying $\alpha=\beta$ within the interval $[0.01,1]$ and $\omega$ within the range $[1,30]$ for the GSS preconditioner.
(b) By varying $\tau$ within the interval $[0.01,0.3]$ and $\bm{\omega}$ within the interval $[1,30]$ for the RGSS-II preconditioner
Figure 2: Convergence curves of the GSS and RGSS-II RGMRES methods varying the parameters $\alpha,$ $\beta,$ $\tau,$ $\bm{\omega}$ for Example 1 with $\bm{\beta}=0.1$.
To demonstrate the influence of the parameter on the performance of the proposed preconditioners, we present graphs of IT counts versus parameters for the GSS and RGSS-II preconditioners in Figure 2. For the GSS preconditioner, we vary $\alpha=\beta$ from $0.01$ to $1$ with a step size of $0.01$ and $\bm{\omega}$ from $1$ to $30$ with a step size of one. For the RGSS-II preconditioner, we vary $\tau$ from $0.01$ to $0.3$ with a step size of $0.01$ and $\bm{\omega}$ from $1$ to $30$ with a step size of one. We can draw the following observation from Figure 2:
  • For both preconditioners, IT exhibits minimal sensitivity to variations in the parameters.
  • Although, for small values of $\bm{\omega}$, IT increases when $\alpha$ increase for GSS preconditioner and $\tau$ increases for RGSS-II preconditioner. Nonetheless, as $\bm{\omega}$ increases, IT decreases, even as the magnitude of $\alpha$ and $\tau$ continue to grow.
Therefore, the proposed preconditioners achieve high efficiency when $\alpha$, $\beta$ and $\tau$ are kept small and $\bm{\omega}$ is large.
Condition number analysis:
Figure 3: Relationship between condition numbers of the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ varying the parameter $\bm{\omega}$ in $[1,100]$ with $\bm{\beta}=0.1$ for Example 1.
To evaluate the robustness of the proposed GSS, RGSS-I and RGSS-II preconditioners, we assess the condition numbers of the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$. For any nonsingular matrix $A$, the condition number $\kappa(A)$ is defined as $\kappa(A):=\|A^{-1}\|_{2}\|A\|_{2},$ where $\|A\|_{2}$ denotes the spectral norm of $A.$ The condition number of $\mathcal{B}$ is $053.6396e+$ 05 which is very large, making the system (1.1) ill-conditioned to solve. In Figure 3, we depict the effect of the parameter $\bm{\omega}$ ranges from 1 to 100 with a step size of one on the preconditioned matrices $\mathcal{M}_{\text{GSS }}^{-1}\mathcal{B},\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ and $\mathcal{M}_{\text{RGSS-I}}^{-1}\mathcal{B}$ for the case def_setup.pow $=5$. We observe that for all values of $\bm{\omega}$, the condition numbers of the preconditioned matrices remain within the range $[1,1.5]$. This indicates that the preconditioned systems are well-conditioned, demonstrating that the GSS, RGSS-I, and RGSS-II preconditioners are robust and effective.
Analyzing the sensitivity of the solution:
Figure 4: Behaviour of the relative errors in the solution with with increasing $n_{p}$ for def_setup.pow $=5,6,7$ with $\bm{\beta}=0.1$ for Example 1.
To assess the reliability of the solution computed using the GSS PGMRES method when the coefficient matrix is slightly perturbed, we consider the perturbed system $(\mathcal{B}+\Delta\mathcal{B})\bm{\widetilde{u}}=\bm{b}$ of the DSPP (1.1). We compute $\Delta\mathcal{B}$ by adding some Gaussian noise to $B$ and $C$ in the following manner:
$\Delta B=\epsilon\cdot n_{p}\cdot std(B).*randn(m,n)\quad\text{and}\quad\Delta C=\epsilon\cdot n_{p}\cdot std(C).*randn(p,m),$
where $n_{p}$ is the noise percentage, $\epsilon=10^{-6},$ $randn(m,n)$ and $std(X)$ denote the random matrices of order $m\times n$ and the standard deviation of the matrix $X,$ respectively. We conducted numerical tests for the GSS PGMRES method for def_setup.pow $=5,6,7$ with $\beta=0.1$. Let $\bm{u}$ and $\bm{\widetilde{u}}$ are the computed solutions of the DSPP and its perturbed system, respectively. We compute the relative error in the solution, given by ${\displaystyle{\frac{\|\bm{\widetilde{u}}-\bm{u}\|_{2}}{\|\bm{u}\|_{2}}}}$ as the noise percentage $n_{p}$ increases from $5\%$ to $40\%$ with step size $5\%.$ These results are presented in Figure 4. We observe that as $n_{p}$ increases, the relative error of the solution consistently remains below $10^{-6}$. This indicates that the solution $\bm{u}$ is insensitive in nature to small perturbations in the coefficient matrix and demonstrates the robustness of the proposed GSS PGMRES method. Similar observations are also noticed for RGSS-I and RGSS-II PGMRES methods.
8 Concluding remarks
This paper introduces three preconditioners, termed GSS, RGSS-I and RGSS-II for solving DSPPs that arise in various applications. We provide a convergence analysis for the GSS iterative method, demonstrating that the method converges for any initial guess vector when the parameter $\bm{\omega}\geq 1/2.$ Moreover, spectral bounds for the preconditioned matrices are derived. Additionally, we have shown that the RGSS-II preconditioner requires almost $m+1$ iterations to solve the DSPP. Numerical experiments for the DSPP arising from the PDE-constrained optimization problem are performed, which demonstrate that the proposed preconditioners are efficient and outperform the existing state-of-the-art preconditioners.
Acknowledgments
Pinki Khatun gratefully acknowledges the Council of Scientific &\&& Industrial Research (CSIR), New Delhi, India, for providing financial support through a fellowship (File No. 09/1022(0098)/2020-EMR-I).